Suppose that your portfolio consists of 20% of XYZ stock and 80% of CCC stock. XYZ has a beta of 1.5 and CCC has a beta of 1.3. What is the expected return of this portfolio according to the CAPM. Assume that the expected return on the market portfolio is 12% and the risk-free rate is 5% Enter your answer as a percent. Do not include the % sign. Round your final answer to two decimals.
Answer #1

Portfolio beta=Respective beta*Respective weight

=(0.2*1.5)+(0.8*1.3)=1.34

Expected return=risk free rate+beta*(market rate-risk free rate)

=5+1.34*(12-5)

=14.38